OpenStudy (anonymous):

3x2(square+5x=2 solve this following

OpenStudy (anonymous):

Is it 3x^2+5x=2

OpenStudy (turingtest):

\[3x^2+5x=2\]???

OpenStudy (anonymous):

yaaaaa

OpenStudy (turingtest):

get it into the form of \[Ax^2+Bx+C=0\]then try to factor it, complete the square, or use the quadratic formula. let me know if/where you get stuck